# A concise tutorial review of Reverse Osmosis and Electrodialysis

### v5

@@ 3.1 Flowrates and rejection in a 1D model of RO for neutral solutes
which is Eq. (30) in ref. [24] and Eq. (34) in ref. [30]. Eq. (21) describes that an RO membrane has a natural limit in what rejection it can achieve, determined by the extent to which solutes are excluded from the membrane (which implies a value of $\Phi_{i}\!<\!1$), and by the extent of solute-membrane friction (which leads to $K_{\text{f},i}\!<\!1$). If a membrane does not do either, i.e., it does not exclude solutes, i.e., $\Phi_{i}\!=\!1$, and it does not impose a frictional force on solutes at all, i.e., $K_{\text{f},i}=1$, then rejection is zero. Thus, for an RO membrane to function, either solutes must be excluded from the membrane, leading to $\Phi_{i}\!<\!1$, or there must be a solute-membrane friction, i.e., $K_{\text{f},i}\!<\!1$, and ideally both.
3.2 A 1D model for RO of neutral solutes including concentration polarization
Next we extend the SF-model with concentration polarization. We can envision this effect as if there is a stagnant layer on the upstream side, in front of the membrane [32, 33]. This layer has a certain thickness and all water and solutes must travel across this layer before reaching the membrane. Eq. (16) which describes solute convection and diffusion in a membrane also applies to the CP layer, now with $\sigma_{i}\!=\!0$, because we are outside the membrane. We then arrive at the exponential law for the CP layer [33]
$c_{\text{int},i}=\left(c_{\text{f},i}-\frac{J_{i}}{v\mathrm{{}_{w}}}\right)\cdot\exp\left(\frac{v\mathrm{{}_{w}}}{k_{\text{cp},i}}\right)+\frac{J_{i}}{v\mathrm{{}_{w}}}$ (22)
where $k_{\text{cp},i}$ is the mass transfer coefficient in the CP layer. With more stirring, or a higher crossflow velocity (flux of the water pumped along the membrane), $k_{\text{cp},i}$ goes up. In Eq. (22) we use ‘f’ for feed, but that is only correct in an experiment at very low water recovery. More generally, when we use this equation in a 2D module calculation, with z a coordinate along the membrane, we should use here the z-dependent concentration in the bulk of the channel. The concentration at the interface between CP layer and membrane (but still in solution) is $c_{\text{int},i}$.
In a 1D geometry where Eq. (15) applies, we can rewrite Eq. (22) to
$c_{\text{int},i}=c_{\text{p},i}+\left(c_{\text{f},i}-c_{\text{p},i}\right)\exp\left(\frac{v\mathrm{{}_{w}}}{k_{\text{cp},i}}\right)$ (23)
and if $c_{\text{int},i}\gg c_{\text{p},i}$, then Eq. (23) simplifies to $c_{\text{int},i}=c_{\text{f},i}\cdot\exp\left(v\mathrm{{}_{w}}/k_{\text{cp},i}\right)$. This result is always valid when the membrane blocks all salts or solutes, because then $c_{\text{p},i}\!=0$. For any asymmetric binary salt, Eqs. (22)–(23) also apply, but the diffusion coefficient $D_{\infty,i}$ is replaced by the harmonic mean diffusion coefficient of the binary salt, $D\mathrm{{}_{hm}}$, see ref. [15].
@@ 3.2 A 1D model for RO of neutral solutes including concentration polarization
$R_{i}=\frac{\left(1-\exp\left(-\text{Pe}_{i}\right)\right)\,\sigma_{i}}{\exp\left(v\mathrm{{}_{w}}/k_{\text{cp},i}\right)\,\left(1-\sigma_{i}\right)+\left(1-\exp\left(-\text{Pe}_{i}\right)\right)\,\sigma_{i}}$ (24)
which is Eq. (13) in ref. [5]. When the mass transfer coefficient of the CP layer is large enough, for instance because of sufficient stirring, then Eq. (24) simplifies to Eq. (18). Eq. (24) predicts a maximum in rejection, as also shown in Fig. 3A, and this maximum rejection is at a membrane Pe-number given by $\text{Pe}_{i}=\ln\left(1+k_{\text{cp},i}/k_{\text{m},i}\right)$ (Eq. (15) in ref. [5]).
@@
Figure 3: Results of calculations for rejection of neutral solutes by an RO membrane in a 1D geometry, combining membrane transport modeling with solute partitioning and concentration polarization, based on Eqs. (18) and (21). A) rejection as function of permeate water flux, $v\mathrm{{}_{w}}$, and solute mass transfer coefficients ($\sigma_{i}\!=\!0.9$). B) Examples of solute concentration profiles across the membrane.
3.3 Pressure and water flow in a 1D model of RO for neutral solutes
Having discussed how solutes are retained by RO membranes, we next describe the relation between the hydrostatic pressure that is applied across an RO membrane, $\Delta P^{\text{h},\infty}$, and the resulting permeate water flux, $v\mathrm{{}_{w}}$. The force balance on the water that will be presented in section 5 describes fluid flow through porous media in the presence of solutes that move through that same porous structure. The theory describes how water flows because of gradients in hydrostatic and osmotic pressure [61] (the latter due to gradients in total solute concentration), and is slowed down by friction with the membrane matrix and by friction with solutes. When solutes also have friction with the membrane (resulting in $K_{\text{f},i}<1$), then the water velocity decreases: with increasing solute-membrane friction, solutes become more of an obstacle for water to flow. In the opposite limit when solutes ‘just flow along’ with the water, without solute-membrane friction, they do not hinder water transport and then for neutral solutes this same theory leads to the classical Darcy equation for fluid flow through a porous medium. At the edges of the membrane, we have jumps both in osmotic pressure and hydrostatic pressure [30, 34, 35], and it is essential to understand these to predict the direction of water flow, especially in the absence of a significant hydrostatic pressure. These topics are addressed in section 5.
@@ 3.3 Pressure and water flow in a 1D model of RO for neutral solutes
The physical phenomena at the outsides of the membrane are highly intriguing, even for neutral solutes. We know there is a jump ‘down’ in solute concentration upon entering the membrane when solutes are excluded from the membrane, described by $\Phi_{i}<1$. A first question is then how the osmotic pressure changes across the interface. In this case also the osmotic pressure goes down, from a value $\Pi^{\infty}$ just outside the membrane to a value $\Pi^{\text{m}}$ just inside the membrane, with $\Pi^{\text{m}}<\Pi^{\infty}$, when $\Phi_{i}\!<\!1$ for all solutes. For ideal solutions the osmotic pressure (both inside and outside a membrane) is proportional to the solute concentration according to $\Pi=c\,RT$ (where c is a total concentration, i.e., a summation over all solutes), and thus the jump in osmotic pressure (counted in the membrane relative to outside) is $\Delta\Pi=\left(c_{\text{m}}-c_{\infty}\right)RT$, where $c_{\infty}$ and $c_{\text{m}}$ are total concentrations of solute. If we now include the partitioning equation, we arrive at $\Delta\Pi=-c_{\infty}\left(1-\Phi_{i}\right)RT$ (assuming that all solutes have the same $\Phi_{i}$), which is negative. This means the osmotic pressure goes down upon entry into the membrane. The magnitude of this change is proportional to the solute concentrations just outside the membrane, and thus $\Delta\Pi$ can be quite significant on the upstream side of a membrane, while on the downstream side the jump in pressure between outside and inside the membrane is much less when the membrane has a high rejection.
The next step is to find the relationship between this osmotic pressure change across the membrane surface and the hydrostatic pressure change there. Because of mechanical equilibrium across this interface, the solvent pressure is invariant across each of the membrane interfaces (between just outside and just inside) [39], and therefore at each of the two membrane interfaces we have
@@
[…]
$v\mathrm{{}_{w}}=\tfrac{1}{RT}\,k_{\text{F-m}}^{\dagger}\,\left(\Delta P^{\text{h},\infty}-\sigma_{i}\cdot\Delta\Pi^{\infty}\right)$ (27)
@@
In Fig. 4A, we analyze how the water flux depends on the pressure. The closer the reflection coefficient, $\sigma_{i}$, is to one, the more clearly there is critical pressure that must be overcome before water starts to flow across the membrane. Furthermore, Fig. 4B illustrates that in the SF model the flow of solutes gradually increases with pressure. This concludes our introduction for the theory of RO for neutral solutes. Advanced models for RO for water desalination (i.e., with ions), in one and two dimensions, are discussed in refs. [3, 33].
Figure 4: Permeate water flux $v\mathrm{{}_{w}}$ (expressed as a Pe-number) and the molar flux, $J_{i}$, of a neutral solute as function of $\Delta P^{\text{h},\infty}$, normalized by the feed osmotic pressure, $c\mathrm{{}_{f}}RT$, across an RO membrane in a 1D geometry. Results of calculations related to Fig. 3 ($k_{\text{cp},i}/k_{\text{m},i}\!=\!5$, $k_{\text{F-m}}^{\dagger}\Pi_{\text{f}}/\left(RTk_{\text{m},i}\right)=1$).
@@ 4 Theory of electrodialysis for a 1:1 salt solution
In this section we describe theory for electrodialysis (ED) in a co-current design, for a 1:1 salt. In ED, feedwater flows into two adjacent sets of channels, see Fig. 1B. In one series of channels the water will be desalinated, which are the diluate channels, denoted by ‘d’. Salt ions move from these d-channels to the concentrate ‘c’ channels. In many practical ED stacks, the direction of flow in a channel is at a ${}^{\circ}$ angle to the flow in adjacent channels. This is called a cross-flow geometry. This geometry is more complicated to model, and thus in this section we describe that we have perfect co-current flow, where the two flows enter the module from the same side and flow through the stack in parallel. The two channels can still be different in width and the flowrates (thus residence times) can be different.
@@
An ED stack consists of many cell pairs, with electrodes on the two sides of the stack, where electronic current becomes ionic current. In this review we do not discuss the electrodes but we focus on the repeating unit of an ED stack, which is the membrane cell pair, see Fig. 1B, which consists of two membranes and two flow channels. As mentioned, in the flow channels we assume co-current flow of the water along the membranes, from inlet to exit of the channel. While the water flows through these thin channels between the membranes, ions move from the d-channels to the c-channels by transport across the membranes. This process is possible because –driven by the electrical current– anions move in one directon out of each d-channel, to be transported through membranes that are selective for anions. These are ‘anion-exchange membranes’ or AEMs, which are membranes in which the water-filled pores are lined with high concentrations of positive membrane charges (of the order of 5 M fixed membrane charge per volume of water in the membrane). The cations move in the other direction, passing cation-exchange membranes, or CEMs. These CEMs have a high concentration of fixed negative charge, again of the order of 5 M, and they preferentially allow access to cations, largely blocking the passage of anions. Each d-channel has such an AEM on one side, and a CEM on the other. The net effect of this layout of a cell pair, of a sequence of an AEM, d-channel, CEM, and c-channel, and this repeated tens to hundreds of times, is that the d-channels are being desalinated, while the salinity in the c-channels increases going from entrance to exit of the channel. Thus in the AEMs anions are the main charge carrier, while in the CEMs cations carry most of the charge. These ions in their respective membrane are the counterions (anions in an AEM, cations in a CEM). The minority ions, that ideally are fully blocked, are called coions (cations in AEM, anions in CEM). Though in ED there will also be water flowing through the membrane, we will neglect that aspect in this tutorial. In this case, water recovery, wr, is directly set by how much of the feedwater flows to the c-channels and how much to the d-channels, namely $\text{{wr}}=1/\left(1+\phi\mathrm{{}_{v,c}}/\phi\mathrm{{}_{v,d}}\right)$. Thus, for equal flow rates to the two channels, we have $\text{{wr}}\!=\!0.50$.
@@
For each channel in the ED cell pair, we set up a steady-state plug flow model, with ion transport through the membranes in a ‘sideways’ direction. In each channel, at each position $z$, where $z$ is the direction along the membranes, there are concentration profiles across the channel, from membrane to membrane, but they can be relatively minor when the channels are narrow. [^2] This is not always the case, and sometimes these profiles matter a lot. Also with three or more different ions, instead of a simple binary salt, the situation in the flow channels can become markedly different, and concentration profiles across the channel must then be modelled. Thus in this tutorial, in each channel we use a z-dependent mass balance that does not have any dependence on position in the perpendicular x-direction, which is the direction towards the membranes, at right angles to the z-direction of water flow along the membranes. Then a mass balance for an ion i at some position in a flow channel (either a d-channnel or c-channel) is
@@
Explanation of coion leakage in ED. Let us describe the functioning of the ED cell pair in an alternative manner. With the current directed to the right, cations will flow rightward, and anions will flow leftward. In the bulk of the channels, for a symmetric system, this current is equally carried by cations, moving right, and anions that move left. Cations and anions can have velocities in this left-right direction that tend to zero (near a blocking membrane), but they do not flow ‘back’ at any point in the module. With AEMs and CEMs alternatingly placed in the stack, and with cations moving right, anions left, then the following will happen. One type of channel has a CEM on its right side. That membrane freely allows the cations to pass to the next channel (to the right). And that same channel has an AEM on its left side, and anions can pass that membrane freely and move left. But now, in these adjacent channels, things are different. The anions want to continue moving left, but now they encounter a CEM which blocks them. And cations that want to move right, encounter an AEM that blocks them. So in this second set of channels, the ions accumulate, and thus the salt concentration increases. These channels will be the concentrate channels. Ions that arrive in these channels, cannot escape. What was just described, is the ideal functioning of an ED stack, and it is generally like this early on in a cell pair, where current densities are still high and concentration differences between the two types of channels, $c\mathrm{{}_{d}}$ and $c\mathrm{{}_{c}}$, still low.1 But further on in the stack, with $c\mathrm{{}_{c}}\gg c\mathrm{{}_{d}}$, ions start to leak from the concentrate channel. They leak in the direction that drives them onward, not back to the diluate channel from which they came, but on to the next diluate channel. The membrane that blocked them from leaving the concentrate channel in the ideal case (e.g., the AEM for a cation), becomes leaky. Why is that? Chemically the membrane is of course the same as before. But this increased coion leakage from the concentrate channels is because the increasing concentration difference across this membrane leads to more and more diffusion of coions. Indeed, diffusion is the main membrane transport mechanism for co-ions in ED. Thus further on in a cell pair, the driving force for diffusion increases, and thus ions that were kept in the concentrate channel quite well at first, now increasingly leak out; so cations that entered a concentrate channel on the left side, now leak out on the right side. This leakage is more pronounced for membranes that have a high diffusion coefficient, and for membranes that are thin. With thin membranes, the gradient in concentration across the membrane is steep, and we therefore have a large driving force for diffusion [41]. To avoid this leakage from the concentrate channel, membranes should not be too thin (at least a thickness of 10 $\mu$m, but 100 $\mu$m is even better). This increased leakage is a reason why, towards the exit of a channel, the net rate of salt removal goes down (per unit membrane area). The other reason is that the local current density goes down, so the transport of ions from diluate to concentrate channels (the intended purpose of ED) also goes down. This decrease is because the resistance goes up in the diluate channel, and Donnan potentials increase. So for a given cell pair voltage, $V\mathrm{{}_{cp}}$, the high current densities that could be achieved early on in the channel, are dropping significantly the further one progresses down the channels. Thus the flux of cations and anions arriving in the concentrate channels goes down significantly. And at the same time diffusional leakage goes up, as discussed. As a consequence, at some point current is still running, but the net salt transport from diluate to concentrate channels (arrival minus leakage) approaches zero. At this point (and beyond), the current efficiency, $\lambda$, is zero. This limiting value of $\lambda=0$ will be used further on to calculate the ‘final’ current density and salt concentrations, $c\mathrm{{}_{d}}$ and $c\mathrm{{}_{c}}$, in a stack run at a fixed cell pair voltage. 1 The same analysis applies to a (short, or ‘differential’) stack that is fed from reservoirs to which the effluent also flows back (this is batch-mode ED operation, where solution is recirculated between the reservoirs and the stack, and is desalinated over time). Then statements about ‘early on in the stack’ can be changed to ‘early in time’, and ‘further on’ by ‘later’, assuming we started with two reservoirs at equal salt concentration. Note that in this layout, the stack can be run at constant voltage or constant current. Instead, in the calculation of co-current flow in a single stack without these reservoirs, then in the common design with two undivided end-electrodes, the entire stack is operated at one value of the cell pair voltage, and the further we are away from the entrance, the lower is the current density.
@@
Eq. (29) introduces a cell-pair based current efficiency $\lambda\mathrm{{}_{cp}}$ which relates to the total salt transport flux between the d- and c-channels, $J\mathrm{{}_{salt}}$, and the current, I, all z-dependent. To calculate $\lambda\mathrm{{}_{cp}}$ and $J\mathrm{{}_{salt}}$ we must solve at each z-coordinate a membrane model for the AEM and the CEM, and in each membrane describe the flow of counterions and coions. In such a model all parameters, including the diffusion coefficients of the ions, and the membrane charge, can be different between AEM and CEM. This calculation is possible, but in this review we simplify this situation, and assume the AEM and CEM are each other’s perfect ‘mirror image’. Thus, the fixed membrane density in an AEM, denoted by the symbol X, has the same magnitude as the membrane charge of a CEM, and only the sign of this membrane charge is opposite. Furthermore, the diffusion coefficient of the counterions in the AEM and CEM is the same, and the same for the coions (coions and counterions can still have diffusion coefficients that are different from one another). With these assumptions, the cell-pair based $\lambda\mathrm{{}_{cp}}$ is now the same as a ‘single membrane’-based current efficiency $\lambda\mathrm{{}_{th}}$, where we use index ‘th’ for ‘theoretical’, because this efficiency cannot be measured but can only be obtained from a theoretical model of a single ion-exchange membrane. Thus, in Eq. (29) we can make the replacement $\lambda\mathrm{{}_{cp}}\rightarrow\lambda\mathrm{{}_{th}}$.
@@
We can combine Eq. (29) for the d- and c-channel, and then integrate between inlet and any position z, which leads to the overall salt mass balance
@@
which expresses that all salt that is removed from one channel ends up in the other channel. This balance assumes there is no water flow across the membranes. If both channels have the same inlet concentration, then $c\mathrm{{}_{d,in}}=c\mathrm{{}_{c,in}}W=c\mathrm{{}_{f}}$. We can now use Eq. (30) to calculated $c\mathrm{{}_{c}}$, and only solve Eq. (29) for the diluate channel. Next, we introduce the time-on-stream, $t^{*}=z/v\mathrm{{}_{z}}$, and rewrite Eq. (29) for the d-channel to

### v6

@@ 3.1 Flowrates and rejection
which is Eq. (30) in ref. [25] and Eq. (34) in ref. [31]. Eq. (21) describes that an RO membrane has a natural limit in what rejection it can achieve, determined by the extent to which solutes are excluded from the membrane (which implies a value of $\Phi_{i}\!<\!1$), and by the extent of solute-membrane friction (which leads to $K_{\text{f},i}\!<\!1$). If a membrane does not do either, i.e., it does not exclude solutes, i.e., $\Phi_{i}\!=\!1$, and it does not impose a frictional force on solutes at all, i.e., $K_{\text{f},i}=1$, then rejection will be zero. Thus, for an RO membrane to function, either solutes must be excluded from the membrane, leading to $\Phi_{i}\!<\!1$, or there must be a solute-membrane friction, i.e., $K_{\text{f},i}\!<\!1$, and ideally both.
3.2 Theory of RO including concentration polarization
Next we extend the SF-model with concentration polarization. We can envision this effect as if there is a stagnant layer on the upstream side, in front of the membrane [33, 34]. This layer has a certain thickness and all water and solutes must travel across this layer before reaching the membrane. Eq. (16), which describes solute convection and diffusion in a membrane, also applies to the CP layer, now with $\sigma_{i}\!=\!0$, because we are outside the membrane. We then arrive at the exponential law for the CP layer [34]
$c_{\text{int},i}=\left(c_{\text{f},i}-\frac{J_{i}}{J\mathrm{{}_{w}}}\right)\cdot\exp\left(\frac{J\mathrm{{}_{w}}}{k_{\text{cp},i}}\right)+\frac{J_{i}}{J\mathrm{{}_{w}}}$ (22)
where $k_{\text{cp},i}$ is the mass transfer coefficient in the CP layer. With more stirring, or a higher crossflow velocity (flux of the water pumped along the membrane), $k_{\text{cp},i}$ goes up. In Eq. (22) we use ‘f’ for feed, but that is only correct in an experiment at very low water recovery. More generally, when we use this equation in a 2D module calculation, with z a coordinate along the membrane, we should use here the z-dependent concentration in the bulk of the channel [4]. The concentration at the interface between CP layer and membrane (but still in solution) is $c_{\text{int},i}$.
In case Eq. (15) applies, we can rewrite Eq. (22) to
$c_{\text{int},i}=c_{\text{p},i}+\left(c_{\text{f},i}-c_{\text{p},i}\right)\exp\left(\frac{J\mathrm{{}_{w}}}{k_{\text{cp},i}}\right)$ (23)
and if $c_{\text{int},i}\gg c_{\text{p},i}$, then Eq. (23) simplifies to $c_{\text{int},i}=c_{\text{f},i}\cdot\exp\left(J\mathrm{{}_{w}}/k_{\text{cp},i}\right)$. This result is always valid when the membrane blocks all salts or solutes, because then $c_{\text{p},i}\!=\!0$. For any asymmetric binary salt, Eqs. (22) and (23) also apply, but the diffusion coefficient $D_{\infty,i}$ is replaced by the harmonic mean diffusion coefficient of the binary salt, $D\mathrm{{}_{hm}}$ [16].
@@ 3.2 Theory of RO including concentration polarization
$R_{i}=\frac{\left(1-\exp\left(-\text{Pe}_{i}\right)\right)\,\sigma_{i}}{\exp\left(J\mathrm{{}_{w}}/k_{\text{cp},i}\right)\,\left(1-\sigma_{i}\right)+\left(1-\exp\left(-\text{Pe}_{i}\right)\right)\,\sigma_{i}}$ (24)
which is Eq. (13) in ref. [5]. Eq. (24) makes use of Eq. (15). When the mass transfer coefficient of the CP layer is large enough, for instance because of sufficient stirring, then Eq. (24) simplifies to Eq. (18). Eq. (24) predicts a maximum in rejection, as also shown in Fig. 3A, and this maximum rejection is at a membrane Pe-number of $\text{Pe}_{i}=\ln\left(1+k_{\text{cp},i}/k_{\text{m},i}\right)$ (Eq. (15) in ref. [5]).
@@
Figure 3: Results of calculations for rejection of neutral solutes by an RO membrane, combining membrane transport modeling with solute partitioning and concentration polarization, based on Eqs. (18) and (21). A) Rejection as function of permeate water flux, $J\mathrm{{}_{w}}$, and solute mass transfer coefficients ($\sigma_{i}\!=\!0.9$). B) Examples of solute concentration profiles across the membrane. All calculations make use of Eq. (15).
3.3 Pressure and water flow in RO for neutral solutes
Having discussed how solutes are retained by RO membranes, we next describe the relation between the hydrostatic pressure that is applied across an RO membrane, $\Delta P^{\text{h},\infty}$, and the resulting permeate water flux, $J\mathrm{{}_{w}}$. The force balance on the water that will be presented in section 5 describes fluid flow through porous media in the presence of solutes that move through that same porous structure. The theory describes how water flows because of gradients in hydrostatic and osmotic pressure [35] (the latter due to gradients in total solute concentration), and is slowed down by friction with the membrane matrix and by friction with solutes. When solutes also have friction with the membrane (resulting in $K_{\text{f},i}<1$), then the water velocity decreases: with increasing solute-membrane friction, solutes become more of an obstacle for water to flow. In the opposite limit when solutes just flow along with the water, without solute-membrane friction, they do not hinder water transport and then for neutral solutes this same theory leads to the classical Darcy equation for fluid flow through a porous medium. At the edges of the membrane, we have jumps both in osmotic pressure and hydrostatic pressure [31, 36, 37], and it is essential to understand these to predict the direction of water flow, especially in the absence of a significant hydrostatic pressure. These topics are addressed in section 5.
@@ 3.3 Pressure and water flow in RO for neutral solutes
For neutral solutes, on the outsides of the membrane, there is a jump down in solute concentration upon entering the membrane when solutes are excluded from the membrane, described by $\Phi_{i}<1$. A first question is then how the osmotic pressure changes across the interface. In this case also the osmotic pressure goes down, from a value $\Pi^{\infty}$ just outside the membrane to a value $\Pi^{\text{m}}$ just inside the membrane, with $\Pi^{\text{m}}<\Pi^{\infty}$, when $\Phi_{i}\!<\!1$ for all solutes. For ideal solutions the osmotic pressure (both inside and outside a membrane) is proportional to the solute concentration according to $\Pi=c\,RT$ (where c is a total concentration, i.e., a summation over all solutes), and thus the jump in osmotic pressure (counted in the membrane relative to outside) is $\Delta\Pi=\left(c_{\text{m}}-c_{\infty}\right)RT$, where $c_{\infty}$ and $c_{\text{m}}$ are total concentrations of solute. If we now include the partition equation, we arrive at $\Delta\Pi=-c_{\infty}\left(1-\Phi_{i}\right)RT$ (assuming that all solutes have the same $\Phi_{i}$), which is negative. This means the osmotic pressure goes down upon entry into the membrane. The magnitude of this change is proportional to the solute concentrations just outside the membrane, and thus $\Delta\Pi$ can be quite significant on the upstream side of a membrane, while on the downstream side the jump in pressure between outside and inside the membrane is much less when the membrane has a high rejection.
The next step is to find the relationship between this osmotic pressure change across the membrane surface and the hydrostatic pressure change there. Because of mechanical equilibrium across this interface, the total pressure is invariant across each of the membrane interfaces (between just outside and just inside) [41], and therefore at each of the two membrane interfaces we have
@@
$J\mathrm{{}_{w}}=-k_{\text{w-m}}^{\dagger}\left(\frac{1}{RT}\frac{\partial P^{\text{h,m}}}{\partial\overline{x}}-\left(1-K_{\text{f},i}\right)\frac{\partial c\mathrm{{}_{m}}}{\partial\overline{x}}\right)$ (26)
where we assume for all solutes the same $K_{\text{f},i}$. The nondimensional coordinate $\overline{x}$ is given by $\overline{x}=x/L\mathrm{{}_{m}}$. Here $k_{\text{w-m}}^{\dagger}$ is a water-membrane permeability that also includes the thickness of the membrane, and indirect friction of water with the membrane via the friction of solutes with the membrane. This latter effect disappears when $K_{\text{f},i}\!=\!1$ and then the $\dagger$-index can be dropped. In that case, i.e., when $K_{\text{f},i}=1$, and if the membrane has constant properties throughout, we have a linear pressure gradient in the membrane that results in water flow (Eq. (65) in ref. [27]). In general, also when $K_{\text{f},i}\!\neq\!1$, when we assume a constant value for $k_{\text{w-m}}^{\dagger}$, we can integrate over the membrane thickness, from $\overline{x}\!=\!0$ to $\overline{x}\!=\!1$, implement Eq. (25) for mechanical equilibrium, as well as the relation for solute partitioning, and then obtain the classical result that the transmembrane water flow rate is given by
$J\mathrm{{}_{w}}=\tfrac{1}{RT}\,k_{\text{w-m}}^{\dagger}\,\left(\Delta P^{\text{h},\infty}-\sigma_{i}\cdot\Delta\Pi^{\infty}\right)$ (27)
@@
In Fig. 4A, we analyze how the water flux depends on the pressure. The closer the reflection coefficient, $\sigma_{i}$, is to one, the more clearly there is a critical pressure that must be overcome before water starts to flow across the membrane. Furthermore, Fig. 4B illustrates that in the SF model the flow of solutes gradually increases with pressure. This concludes our introduction for the theory of RO for neutral solutes. Advanced models for RO for water desalination (i.e., with ions), in one and two dimensions, are discussed in refs. [3, 34, 14]. It is also explained there how the osmotic pressure in a salt solution is lower than the ideal pressure because of Coulombic interactions between ions, as was addressed in section 2.3.2.
Figure 4: Permeate water flux $J\mathrm{{}_{w}}$ (expressed as a Pe-number) and the molar flux of neutral solutes, $J_{i}$, as function of $\Delta P^{\text{h},\infty}$, normalized by the feed osmotic pressure, $c\mathrm{{}_{f}}RT$, across an RO membrane. Results of calculations related to Fig. 3 ($k_{\text{cp},i}/k_{\text{m},i}\!=\!5$, $k_{\text{w-m}}^{\dagger}\Pi_{\text{f}}/RTk_{\text{m},i}=1$).
@@ 4 Theory of electrodialysis for a 1:1 salt solution
In this section we describe theory for electrodialysis (ED) in a co-current design, for a 1:1 salt. In ED, feedwater flows into two adjacent sets of channels, see Fig. 1B. In one series of channels the water will be desalinated, which are the diluate channels, denoted by ‘d’. Salt ions move from these d-channels to the concentrate ‘c’ channels. In many practical ED stacks, the direction of flow in a channel is at a 90 angle to the flow in adjacent channels. This is called a cross-flow geometry. This geometry is more complicated to model, and thus in this section we describe that we have perfect co-current flow, where the two flows enter the module from the same side and flow through the stack in parallel. The two channels can still be different in width and the flowrates (thus residence times) can be different.
@@
An ED stack consists of many cell pairs, with electrodes on the two sides of the stack, where electronic current becomes ionic current. In this review we do not discuss the electrodes but we focus on the repeating unit of an ED stack, which is the membrane cell pair, see Fig. 1B, which consists of two membranes and two flow channels. As mentioned, in the flow channels we assume co-current flow of the water along the membranes, from inlet to exit of the channel. While the water flows through these thin channels between the membranes, ions move from the d-channels to the c-channels by transport across the membranes. This process is possible because –driven by the electrical current– anions move in one direction out of each d-channel, to be transported through membranes that are selective for anions. These are anion-exchange membranes (AEMs), which are membranes in which the water-filled pores are lined with high concentrations of positive membrane charges (of the order of 5 M fixed membrane charge per volume of water in the membrane). The cations move in the other direction, passing cation-exchange membranes (CEMs). These CEMs have a high concentration of fixed negative charge, again of the order of 5 M, and they preferentially allow access to cations, largely blocking the passage of anions. Each d-channel has such an AEM on one side, and a CEM on the other. The net effect of this layout of a cell pair, of a sequence of an AEM, d-channel, CEM, and c-channel, and this repeated tens to hundreds of times, is that the d-channels are being desalinated, while the salinity in the c-channels increases going from entrance to exit of the channel. Thus in the AEMs anions are the main charge carrier, while in the CEMs cations carry most of the charge. These ions in their respective membrane are the counterions (anions in an AEM, cations in a CEM). The minority ions, that ideally are fully blocked, are called coions (cations in AEM, anions in CEM). Though in ED there is also water flowing through the membrane, we will neglect that aspect in this tutorial. Water recovery, wr, is then directly set by how much of the feedwater flows to the c-channels and how much to the d-channels, namely $\text{{wr}}=1/\left(1+\phi\mathrm{{}_{v,c}}/\phi\mathrm{{}_{v,d}}\right)$. Thus, for equal flow rates to the two channels, we have $\text{{wr}}\!=\!0.50$.
@@
For each channel in the ED cell pair, we set up a steady-state plug flow model, to describe the flow from inlet to exit of a channel, with ion transport through the membranes included, directed sideways. In each channel, at each position $z$, where $z$ is the direction along the membranes, there are concentration profiles across the channel, from membrane to membrane, but they can be relatively minor when the channels are narrow. [^2] This is not always the case, and sometimes these profiles matter a lot. Also with three or more different ions, instead of a simple binary salt, the situation in the flow channels can be markedly different, and concentration profiles across the channel must then be modelled. Thus in this tutorial, in each channel we use a z-dependent mass balance that does not have any dependence on position in the perpendicular x-direction, which is the direction towards the membranes, at right angles to the z-direction of water flow along the membranes. Then a mass balance for an ion i at some position in a flow channel (either a d-channnel or c-channel) is
@@
Eq. (29) introduces a cell-pair based current efficiency $\lambda\mathrm{{}_{cp}}$ which relates to the total salt transport flux between the d- and c-channels, $J\mathrm{{}_{salt}}$, and the current, I, all z-dependent. To calculate $\lambda\mathrm{{}_{cp}}$ and $J\mathrm{{}_{salt}}$ we must solve at each z-coordinate a membrane model for the AEM and the CEM, and in each membrane describe the flow of counterions and coions. In such a model all parameters, including the diffusion coefficients of the ions, and the membrane charge, can be different between AEM and CEM. This calculation is possible, but in this review we simplify this situation, and assume the AEM and CEM are each other’s perfect mirror image. Thus, the fixed membrane density in an AEM, denoted by the symbol X, has the same magnitude as the membrane charge of a CEM, and only the sign of this membrane charge is opposite. Furthermore, the diffusion coefficient of the counterions in the AEM and CEM is the same, and the same for the coions (coions and counterions can still have diffusion coefficients that are different from one another). With these assumptions, the cell-pair based $\lambda\mathrm{{}_{cp}}$ is now the same as a single membrane-based current efficiency $\lambda\mathrm{{}_{th}}$, where we use index ‘th’ for theoretical, because this efficiency cannot be measured but can only be obtained from a theoretical model of a single ion-exchange membrane. Thus, in Eq. (29) we make the replacement $\lambda\mathrm{{}_{cp}}\rightarrow\lambda\mathrm{{}_{th}}$.
@@
We combine Eq. (29) for the d- and c-channel, and then integrate between inlet and any position z, which leads to the overall salt mass balance
@@
which expresses that all salt that is removed from one channel ends up in the other channel. This balance assumes there is no water flow across the membranes. If both channels have the same inlet concentration, then $c\mathrm{{}_{d,in}}=c\mathrm{{}_{c,in}}=c\mathrm{{}_{f}}$. We can now use Eq. (30) to calculated $c\mathrm{{}_{c}}$, and only solve Eq. (29) for the diluate channel. Next, we introduce the time-on-stream, $t^{*}=z/v\mathrm{{}_{z}}$, and rewrite Eq. (29) for the d-channel to